OpenStudy (anonymous):

can someone help me with evaluate the expressions (-4/7)^0 and -(5)^0

OpenStudy (anonymous):

any number to the power of zero is ONE so the first one is ONE

OpenStudy (anonymous):

the second on is not one, because the minus sign is out front that makes the second one MINUS ONE

OpenStudy (anonymous):

just 1 that's it? so the answer is 1? and how is that possible ?

OpenStudy (anonymous):

how is what possible?

OpenStudy (anonymous):

im sorry im so dumb at math im trying to learn im just wondering how does (-4/7)^0 turn into 1

OpenStudy (anonymous):

\[b^0=1\] always

OpenStudy (danjs):

For example \[\frac{ 3^2 }{ 3^2 } = 1 = \frac{ 3^{2-2} }{ 1 } = 3^0\]

OpenStudy (anonymous):

it is just notation that is all don't try to think too hard about it like "multiply a number by itself zero times" means nothing

OpenStudy (anonymous):

it makes the rules of exponents consistent is all, so it is defined that way

OpenStudy (anonymous):

ok can you help me with more?

OpenStudy (danjs):

A number divided by itself equals 1, following the exponent rules, you can also say a number to zero power is 1.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

can you help me with more?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

evaluate 10^-6 is it -60?

OpenStudy (anonymous):

no it is certainly not

OpenStudy (anonymous):

\[\huge b^{-n}=\frac{1}{b^n}\] \[\huge 10^{-6}=\frac{1}{10^6}\]

OpenStudy (anonymous):

\[10^6=1,000,000\] so \[10^{-6}=\frac{1}{1,000,000}\]

OpenStudy (anonymous):

so 1/10^6 is the answer? im trying to figure out how to solve it because my books make no since to me

OpenStudy (anonymous):

yes, the answer is \(\frac{1}{10^6}\) or \(\frac{1}{1,000,000}\) or \(0.000001\)

OpenStudy (anonymous):

ok what about rewrite the expression without using a negative exponent 1/6x^-5 simplify as much as possible

OpenStudy (anonymous):

the minus sign in the exponent means the reciprocal if it is up, bring it down if it is down, bring it up

OpenStudy (anonymous):

\[\frac{1}{6x^{-5}}=\frac{x^5}{6}\]

OpenStudy (anonymous):

your awesome! finally something makes since lol I have a few more questions if you don't mind?

OpenStudy (anonymous):

go ahead although i have to run soon

OpenStudy (anonymous):

write p^2 (p^5) without exponents and fill in the blank p^2 (p5) =p?

OpenStudy (anonymous):

without exponents?

OpenStudy (anonymous):

it would just be \[ppppppp\] seven copies of \(p\)

OpenStudy (anonymous):

\[p^2\times p^5=p^{2+5}=p^7\] easier to write with exponents

OpenStudy (anonymous):

ok thank you only a few more im sorry for bothering you

OpenStudy (anonymous):

x(x^2)(x^2) simplify

OpenStudy (anonymous):

\[x^{1+2+2}=x^5\]

OpenStudy (anonymous):

multiply 5v(3v^5w^3)2w^8 simplify as much as possible

OpenStudy (anonymous):

\[5v(3v^5w^3)2w^8\] not sure what the parentheses are for

OpenStudy (anonymous):

multiply the numbers, you get \(30\)

OpenStudy (anonymous):

I just put them there because I didn't know what else to put for multiplication

OpenStudy (anonymous):

add the exponents you get \[\large 30v^6w^{11}\]

OpenStudy (anonymous):

simplify v^-3 (v^-9) write only with positive exponents only